# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ 7. Test Control of CAT
When implementing a CAT system, in addition to considering the three components mentioned above, several key factors need to be taken into consideration. These factors include content control, fairness, robustness, and search efficiency.
7.1. Content Control
@@
Content and exposure control is a secondary, but crucial, goal in CAT. It aims to balance the frequency of each question’s use from the question bank. Proper content control can help mitigate the risk of overexposure of questions, minimize question waste, and maximize test coverage. Two popular strategies for exposure control are the Sympson-Hetter method (Sympson and Hetter, 1985) and the A-Stratified method (Chang and Ying, 1999): (1) Sympson-Hetter Method manages question exposure rates using conditional probabilities. It doesn’t assign a selected question to the examinee immediately; instead, it passes through a probability filter. The actual chance a question is given to an examinee depends on both its selection likelihood and a content control parameter. Adjustments to the parameter aim to keep question exposure within acceptable limits. However, this method may not effectively increase the usage rate of low-exposure questions and doesn’t consider the match between question difficulty and examinee proficiency. Enhancements to this method have been developed to address these limitations (van der Linden and Veldkamp, 2004, 2007; Barrada et al., 2009b); (2) A-Stratified Method and its subsequent researches (Chang et al., 2001; Barrada et al., 2006, 2014) are designed to counteract selection biases of algorithms favoring certain questions (e.g., Fisher Information prefers highly differentiated questions). It first stratifies the question bank by difficulty, and starts with less discriminating sub-banks and progresses to more challenging ones as the test unfolds. This stratification ensures a more even exposure of all questions. On the other hand, numerous studies (Bi et al., 2020; Wang et al., 2023a) have attempted to incorporate the coverage of knowledge concepts as a criterion in question selection, aiming to make the assessment more comprehensive.
7.2. Fairness
@@
Fairness is a topic of profound societal significance in both education and machine learning research fields, sparking numerous discussions and leading to the development of many fairness-aware learning algorithms (Cleary, 1968; Chai and Wang, 2022; Li and Liu, 2022). As a technology with potential applications in high-stakes testing, fairness in CAT is a paramount concern. In CAT, the bias that leads to fairness issues can be introduced through three components: the Cognitive Diagnosis Model (CDM), the selection algorithm, and the question bank:
@@
  • Bias in CDMs. Biases in CDMs may stem from the skewed training data, which could reflect the underrepresentation of certain groups or pre-existing educational disparities (Liu et al., 2022; Thompson, [n.d.]; Kizilcec and Lee, 2022). Such biases can lead to an inaccurate and biased estimation of an examinee’s proficiency $\theta$, resulting in unfair outcomes.
@@
  • Bias in Question Bank. The question bank may contain biases if questions are not equally applicable or relatable to all examinees, potentially disadvantaging certain groups (Camilli and Shepard, 1994; Hambleton et al., 1991). For example, some questions in NAPLAN have been deemed unfair for rural examinees, as these questions don’t relate to their real-life experiences (Roberts, [n.d.]). Various methods have been proposed to detect this type of bias (Chu and Lai, 2013; Camilli and Shepard, 1994; Mellenbergh, 1989).
@@
  • Bias in Selection Algorithms. Selection algorithm can introduce bias since every algorithm has its own “selection preferences”. For example, the Maximum Fisher Information tends to select questions with high discrimination (Lord, 2012). If such questions unexpectedly correlate with specialized knowledge known only to a specific group, bias may ensue.
Figure 10. Example of a test-taker’s GRE score report. The GRE General Test provides three scaled scores: Verbal Reasoning, Quantitative Reasoning, and Analytical Writing. The percentile scores facilitate comparison with the performance of other test-takers within a specified comparison group.
@@
Concerns about fairness in CAT also stem from the fact that examinees answer different questions (Green et al., 1984). Equating, a technique used to ensure score equivalence across different tests, is commonly employed to address such concerns (Brigman and Bashaw, 1976). Many further studies about equating scores have been conducted (van der Linden, 2000; JAN, [n.d.]; Sawaki, 2001). Equating ensures fairness by satisfying three requirements (van der Linden, 2000; Braun, 1982): (1) Measurement of the Same Variable: the proficiency estimated by the CAT must reflect the same aspects of proficiency; (2) Equity: an examinee’s score on one test after equating should match the score they would have received if they had taken this test directly; (3) Invariance across Populations: every examinee should get the same equated score for the same performance in different populations. In real-world tests such as the GRE, equating has been used to standardize scores and percentiles (Figure 10), taking into account the difficulty of the questions answered. This process ensures that scores can be compared fairly across different examinees worldwide.
7.3. Robustness
@@
Noise in CAT can impact the precision of the estimated proficiency of an examinee, leading to potential errors in score interpretation. In CAT, noise usually refers to the random variability or measurement error that can affect the accuracy of estimation. It can arise from various sources such as test administration conditions, test-taker behavior, or question characteristics. To mitigate the effects of noise in CAT, a robustness factor is introduced to help stabilize the estimation of proficiency by incorporating additional information, thereby counteracting the impact of noise and improving the reliability (Baker and Kim, 2004). In machine learning, various robustness techniques are employed to enhance the performance of models in the presence of noise, such as regularization methods (Li and Zhang, 2021), data augmentation (Wang et al., 2021), adversarial methods (Dong et al., 2020), ensemble methods (Kariyappa and Qureshi, 2019). In the CAT testing process, significant sources of noise such as guess and slip factors made by examinees, introduce uncertainty. For example, an examinee’s proficiency level may not be uniquely determined by their responses, as they may solve a particular question correctly using different knowledge concepts or even by guessing. The presence of this noise and uncertainty poses a significant challenge to the robustness of CAT. Veldkamp and Verschoor (2019) consider the uncertainty in the question parameters in the selection process. Recently, Zhuang et al. (2022a) explores the use of ensemble learning to combine multiple potential estimates at each step to enhance the proficiency estimation.
7.4. Search Efficiency
In large-scale educational testing, such as the GRE (Mills and Steffen, 2000), the search efficiency of selection algorithms poses a significant challenge. The algorithms typically evaluate the value of each question in a brute-force manner. Consequently, the time complexity of selecting a question in each testing step is linear to the number of questions in the bank $\mathcal{Q}$, denoted as $O(|\mathcal{Q}|)$. This linear relationship is a bottleneck in intelligent educational systems. Many organizations like GMAT (Rudner, 2009) resort to the manual development of filtering rules by experts, a labor-intensive process that somewhat limits the efficacy of CAT. To address the efficiency issue in question selection, researchers have proposed two kinds of solutions: (1) The application of the Particle Swarm Optimization (PSO) algorithm in IRT has proven effective in mitigating computational complexities (Huang et al., 2009). PSO operates with particles concurrently exploring the search space, where each particle represents a potential solution to the optimization problem of question selection. This parallel exploration allows PSO to cover a larger portion of the search space in each iteration, improving the chances of rapidly identifying optimal questions; (2) Leveraging efficient search methods from the realms of recommendation systems and information retrieval to enhance CAT’s search efficiency, e.g., creating of a tree-based index (Ram and Gray, 2012; Zhu et al., 2018). This approach organizes data in hierarchical tree structures, with balanced trees ensuring logarithmic time complexity for search operations. Recently, Hong et al. (2023) propose a Search-Efficient CAT framework, which employs examinee-aware space partitioning to construct a tree-based index. It reduces the search space and eliminates redundant searches in multi-round testing steps, thereby reducing the search complexity from $O(|\mathcal{Q}|)$ to $O(\log|\mathcal{Q}|)$.
Discussion 7.
@@ 7.4. Search Efficiency
Table 2 showcases the underlying causes and advantages of different factors in CAT test control. While accuracy and efficiency stand as primary objectives, these factors hold significant importance for practical settings, especially in high-stakes testing scenarios (e.g., competitive or selective examinations). However, consideration of these factors may inevitably reduce accuracy. For example, when considering the additional fairness to ensure equity among different groups, it might be necessary to deviate from the optimal trajectory of a well-trained selection algorithm. Thus, CAT poses a multidimensional decision-making challenge, necessitating the consideration of various factors at the same time using diverse machine learning techniques.
Table 2. Test Control: Key Factors in CAT Implementation
Factors Category Causes Advantages Pubs
Content Control
Exposure
Control
Unbalanced question usage
Mitigates overexposure;
Test security
(Sympson and Hetter, 1985; Chang and Ying, 1999)
(Barrada et al., 2009b, 2014)
Test Coverage Algorithmic preferences
Comprehensive assessment;
Broad knowledge coverage
(Bi et al., 2020; Wang et al., 2023a)
Fairness Bias in CDMs
Skewed training data;
Underrepresentation of certain groups
Promotes equitable outcomes;
Improves accuracy of proficiency estimation
(Kizilcec and Lee, 2022; Thompson, [n.d.])
(Liu et al., 2022)
Bias in Question
Bank
Unequal applicability;
Cultural or regional biases
Ensures content relevance;
Reduces disadvantage for certain groups
(Camilli and Shepard, 1994; Hambleton et al., 1991)
(Roberts, [n.d.]; Chu and Lai, 2013)
Bias in Selection
Algorithms
Algorithmic preferences
Reduces disadvantage for certain groups
(Lord, 2012)
Equating
Different selected questions across examinees
Score comparability;
Fairness across different tests
(Green et al., 1984; van der Linden, 2000)
(JAN, [n.d.]; Sawaki, 2001)
Robustness Noise Resistance
Random variability;
Guessing and slipping factors
Stabilizes estimation;
Improves reliability
(Baker and Kim, 2004; Zhuang et al., 2022a)
Modeling Uncertainty
Uncertainty in response
Improves accuracy of proficiency estimation
(Veldkamp and Verschoor, 2019)
(Zhuang et al., 2022a)
Search Efficiency
Large question banks;
Brute-force search
Reduces search complexity (Huang et al., 2009; Hong et al., 2023)
8. Evaluation
@@
CAT aims to estimate examinees’ proficiencies with high accuracy and efficiency. While the true proficiency is inherently unobservable and cannot serve as a direct benchmark, various metrics have been developed to assess the performance of CAT methods, such as correlation coefficients, bias, and measurement error (Crocker and Algina, 1986; Van der Linden and Glas, 2010). This section introduces two of the most extensively utilized evaluation methods: simulation of proficiency estimation and examinee score prediction.
Simulation of Proficiency Estimation.
@@
The simulation of ability estimation is a foundational evaluation technique in CAT (Vie et al., 2017). Since we don’t have direct access to an examinee’s true proficiency level $\theta_{0}$, we create a simulated version of it. By sampling a set of true proficiency values $\{\theta_{0}^{1},\theta_{0}^{2},...,\theta_{0}^{N}\}$ from the distribution of examinee proficiencies, we construct a virtual cohort of $N$ ”examinees”. This approach enables us to further emulate the interactions between examinees (with these proficiencies) and any question from the question bank, utilizing CDMs. Consequently, upon the conclusion of a CAT assessment, the estimated proficiency values $\hat{\theta}^{T}$ can be directly compared with the true values $\theta_{0}$. For instance, by computing the Mean Square Error (MSE) between these sets of values $\mathbb{E}\|\hat{\theta}^{t}-\theta_{0}\|$, we gain an accurate assessment of the CAT system’s precision in estimating examinee proficiencies (Bi et al., 2020; Cheng, 2009).
Examinee Score Prediction.
@@
In CAT systems powered by machine learning, proficiency estimates are validated by predicting examinee responses (correct or incorrect) on a held-out test set. For example, we can allocate 70%-20%-10% examinees for training, validation, and testing respectively, and the examinees in the validation/testing set won’t appear in training. Training data calibrates question parameters (Sections 6.1) and optimizes data-driven selection algorithms (Sections 5.3, 5.4). During validation or testing, the responses of each examinee are further divided into a candidate set $\mathcal{Q}_{i}$ for selecting questions and a held-out meta set $\mathcal{M}_{i}$ for evaluation. The candidate set $\mathcal{Q}_{i}$ (with corresponding response label $y$) is used to simulate the CAT procedure: It involves selecting questions from $\mathcal{Q}_{i}$, updating proficiency estimates after each step, and then accessing estimate’s precision by predicting responses on $\mathcal{M}_{i}$. The assumption is that better score predictions reflect more accurate proficiency estimates. Thus, the binary classification metrics can be used for evaluations, e.g., Prediction Accuracy (ACC) (Gao et al., 2021) and Area Under ROC Curve (AUC) (Bradley, 1997).
8.1. Datasets
@@
To rigorously assess the efficacy of a CAT solution, it is essential to deploy diverse datasets that can effectively challenge and evaluate the system’s capabilities. Three types of data can be employed for this evaluation: (1) Educational Assessment Datasets: These datasets originate from actual standardized tests such as the SAT, GRE, or GMAT. They encompass a broad spectrum of questions covering multiple subjects, coupled with responses from a substantial pool of test-takers. Such datasets are invaluable for gauging the CAT system’s capacity. However, due to stringent privacy laws, intellectual property rights, and the protective measures taken by educational testing companies regarding their questions and datasets, most of these are proprietary and not easily accessible. (2) Real-World Educational Data: This category includes data collected from educational environments in practice, such as schools, universities, and online learning platforms. It provides insights into how examinees interact with educational content and assessments in a natural setting. The data may encompass examinee information, performance responses, learning behaviors, question characteristics, etc. We have open-sourced a comprehensive education-related dataset library: https://github.com/bigdata-ustc/EduData. It includes a range of publicly available datasets along with previously private datasets. Additionally, we have provided a detailed data analysis to support further CAT research and application in educational settings. The following are introductions to several commonly used datasets, and more datasets can be found at the EduData gitHub link.
  • ASSISTments (Feng et al., 2009), established in 2004, is an online tutoring platform in the United States that offers examinees both assessments and instructional support. To date, the ASSISTments team has released four public datasets [^2] https://sites.google.com/site/assistmentsdata/datasets/: ASSISTments2009, ASSISTments2012, ASSISTments2015, and ASSISTments2017. These datasets are response data and mostly collected from mathematics in middle school. They also include valuable side information, such as attempt count (the number of tries an examinee has made), ms first response (the time it takes for an examinee’s first response), problem type, and average confidence.
  • Junyi Dataset (Chang et al., 2015) includes logs and exercise data from Junyi Academy, a Chinese online learning platform launched in 2012 using Khan Academy’s open-source code. It features a detailed question hierarchy and relationships, labeled by experts.
  • MOOCCube [^3] https://www.biendata.xyz/competition/chaindream_mooccube_task2/, Massive Open Online Courses (MOOCs) are among the most prevalent platforms for online learning. This dataset collects examinees’ responses to questions related to various computer science knowledge concepts. Additionally, the dataset includes the text of the problems, which can be used to enhance the performance of question selection, proficiency estimation, question characteristics analysis, etc.
  • EdNet Dataset (Choi et al., 2020) is a large collection of examinee learning records from the AI tutoring system Santa [^4] https://github.com/riiid/ednet, which is used for English language learning in South Korea. It focuses on examinees preparing for the eTOEIC (Test of English for International Communication) Listening and Reading Test, with over 131 million learning records from approximately 784,000 examinees.
  • Eedi2020 Dataset (Wang et al., 2020a), released for the NeurIPS 2020 Education Challenge, contains over 17 million records of examinees’ responses to mathematics multiple-choice questions on the Eedi platform [^5] https://eedi.com/projects/neurips-education-challenge. It includes detailed information on examinees’ choices, demographics, and containment relationships of knowledge concepts, as well as associated quiz and curriculum metadata. This extensive dataset enables in-depth analysis of examinee behaviors and the development of personalized tools.
@@
(3) Simulated Datasets. These are artificially created datasets that mimic the characteristics of real examinee responses as illustrated above. They can be tailored to include specific patterns, noise levels, and distributions, allowing for controlled testing of the CAT system under various scenarios. Monte Carlo simulations can also be used to generate datasets with known properties and ground truth (Van der Linden and Glas, 2010). These datasets are useful for validating the CAT system’s capability to estimate proficiencies accurately and to adapt to the simulated changes during the testing process.
Table 3. AUC results reported by different CAT methods
ASSISTments Eedi2020
Selection Algorithm CDM IRT (Embretson and Reise, 2013) NeuralCD (Wang et al., 2020c) IRT (Embretson and Reise, 2013) NeuralCD (Wang et al., 2020c)
AUC@5 AUC@20 AUC@5 AUC@20 AUC@5 AUC@20 AUC@5 AUC@20
Random
67.68 (Wang et al., 2023a)
70.68 (Zhuang et al., 2023)
65.86 (Ma et al., 2023)
68.43 (Wang et al., 2023a)
72.61 (Zhuang et al., 2023)
67.73 (Wang et al., 2023a)
71.19 (Zhuang et al., 2023)
70.52 (Yu et al., 2023)
69.70 (Wang et al., 2023a)
72.83 (Zhuang et al., 2023)
68.38 (Wang et al., 2023a)
69.05 (Zhuang et al., 2023)
71.98 (Wang et al., 2023a)
74.82 (Zhuang et al., 2023)
68.45 (Wang et al., 2023a)
69.32 (Zhuang et al., 2023)
73.67 (Yu et al., 2023)
72.98 (Wang et al., 2023a)
74.99 (Zhuang et al., 2023)
Statistical Algorithms Fisher Information (Lord, 2012)
67.95 (Wang et al., 2023a)
71.33 (Zhuang et al., 2023)
66.41 (Ma et al., 2023)
69.26 (Wang et al., 2023a)
73.54 (Zhuang et al., 2023)
68.92 (Wang et al., 2023a)
70.60 (Zhuang et al., 2023)
72.66 (Wang et al., 2023a)
76.24 (Zhuang et al., 2023)
KL Information (Chang and Ying, 1996)
67.92 (Wang et al., 2023a)
71.38 (Zhuang et al., 2023)
69.23 (Wang et al., 2023a)
73.57 (Zhuang et al., 2023)
68.69 (Wang et al., 2023a)
69.79 (Zhuang et al., 2023)
72.60 (Wang et al., 2023a)
75.73 (Zhuang et al., 2023)
Fisher Information
+ Robust (Zhuang et al., 2022a)
68.93 (Zhuang et al., 2022a) 75.99 (Zhuang et al., 2022a)
KL Information
+ Robust (Zhuang et al., 2022a)
68.90 (Zhuang et al., 2022a) 76.03 (Zhuang et al., 2022a)
Active Learning MAAT (Bi et al., 2020)
68.24 (Wang et al., 2023a)
71.54 (Zhuang et al., 2023)
66.24 (Ma et al., 2023)
69.7 (Wang et al., 2023a)
73.08 (Zhuang et al., 2023)
67.96 (Wang et al., 2023a)
70.98 (Zhuang et al., 2023)
70.85 (Yu et al., 2023)
71.17 (Wang et al., 2023a)
72.27 (Zhuang et al., 2023)
69.09 (Wang et al., 2023a)
70.32 (Zhuang et al., 2023)
73.19 (Wang et al., 2023a)
74.46 (Zhuang et al., 2023)
69.03 (Wang et al., 2023a)
70.12 (Zhuang et al., 2023)
74.33 (Yu et al., 2023)
73.75 (Wang et al., 2023a)
75.83 (Zhuang et al., 2023)
MAAT
+ Robust (Zhuang et al., 2022a)
68.93 (Zhuang et al., 2022a) 76.09 (Zhuang et al., 2022a) 70.39 (Zhuang et al., 2022a) 76.63 (Zhuang et al., 2022a)
Reinforcement Learning GMOCAT (Wang et al., 2023a) 69.13 (Wang et al., 2023a) 71.91 (Wang et al., 2023a) 69.95 (Wang et al., 2023a) 72.95 (Wang et al., 2023a) 69.81 (Wang et al., 2023a) 74.19 (Wang et al., 2023a) 71.25 (Wang et al., 2023a) 75.76 (Wang et al., 2023a)
NCAT (Zhuang et al., 2022b)
68.67 (Wang et al., 2023a)
71.53 (Zhuang et al., 2023)
71.06 (Wang et al., 2023a)
73.50 (Zhuang et al., 2023)
69.28 (Wang et al., 2023a)
71.59 (Zhuang et al., 2023)
72.53 (Yu et al., 2023)
71.68 (Wang et al., 2023a)
73.59 (Zhuang et al., 2023)
69.04 (Wang et al., 2023a)
72.11 (Zhuang et al., 2023)
73.32 (Wang et al., 2023a)
76.66 (Zhuang et al., 2023)
69.09 (Wang et al., 2023a)
74.10 (Zhuang et al., 2023)
74.49 (Yu et al., 2023)
74.55 (Wang et al., 2023a)
79.12 (Zhuang et al., 2023)
Meta Learning Algorithms BOBCAT (Ghosh and Lan, 2021)
68.65 (Wang et al., 2023a)
71.68 (Zhuang et al., 2023)
66.41 (Ma et al., 2023)
70.97 (Wang et al., 2023a)
73.39 (Zhuang et al., 2023)
69.50 (Wang et al., 2023a)
71.45 (Zhuang et al., 2023)
71.98 (Yu et al., 2023)
71.80 (Wang et al., 2023a)
72.84 (Zhuang et al., 2023)
68.94 (Wang et al., 2023a)
74.42 (Zhuang et al., 2023)
73.24 (Wang et al., 2023a)
76.58 (Zhuang et al., 2023)
69.17 (Wang et al., 2023a)
76.00 (Zhuang et al., 2023)
75.12 (Yu et al., 2023)
74.51 (Wang et al., 2023a)
79.00 (Zhuang et al., 2023)
DL-CAT (Ma et al., 2023) 66.68 (Ma et al., 2023)
SACAT (Yu et al., 2023) 75.24 (Yu et al., 2023) 75.48 (Yu et al., 2023)
Subset Selection
Algorithms
BECAT (Zhuang et al., 2023) 71.44 (Zhuang et al., 2023) 73.61 (Zhuang et al., 2023) 71.60 (Zhuang et al., 2023) 73.70 (Zhuang et al., 2023) 73.15 (Zhuang et al., 2023) 76.82 (Zhuang et al., 2023) 76.30 (Zhuang et al., 2023) 79.36 (Zhuang et al., 2023)
Table 3 displays some representative methods of each category of selection algorithms and their AUC results on two different datasets. The comparison in this survey focuses on the results at the early testing stage (step=5) and the final testing stage (step=20). It is important to note that the results cannot be directly compared if experimental settings are not standardized. Despite this, the table as a whole reveals that data-driven methods (e.g., reinforcement learning, meta-learning methods) generally outperform statistical methods. This is because these methods can train and optimize selection algorithms from examinee large-scale response data, while statistical methods simply adhere to fixed functions for selecting questions. The latest subset selection methods do not require training but are remarkably effective. This is primarily because they attempt to explicitly approximate the objectives of CAT and provide theoretical guarantees on estimation errors. Furthermore, it is observed that considering factors within test control, such as robustness, can enhance accuracy.
9. Opportunities for Future Research

### 09 Mar 2026 v3

@@ 7 Test Control of CAT
When implementing a testing system, in addition to considering the three components mentioned above, several key factors need to be taken into consideration, such as exposure control, fairness, robustness, and search efficiency.
7.1 Exposure Control
@@
Exposure control aim to balance the frequency of each question’s use from the question bank. Proper exposure control can help mitigate the risk of overexposure of questions, minimize question waste, and maximize test coverage. Two popular strategies for exposure control are the Sympson-Hetter method [150] and the A-Stratified method [24]: (1) Sympson-Hetter Method manages question exposure rates using conditional probabilities. It doesn’t assign a selected question to the examinee immediately; instead, it passes through a probability filter. The actual chance a question is given to an examinee depends on both its selection likelihood and a exposure control parameter, keeping question exposure within acceptable limits. However, this method may not effectively increase the usage rate of low-exposure questions. Enhancements to this method have been developed to address these limitations [156, 157, 12]; (2) A-Stratified Method and its subsequent researches [22, 11, 9] are designed to counteract selection biases of algorithms favoring certain questions (e.g., Fisher Information prefers highly differentiated questions). On the other hand, numerous studies [16, 168, 122] have attempted to incorporate the coverage of knowledge concepts as a criterion in question selection, aiming to make the assessment more comprehensive.
This factor is crucial for both humans and AI. When students are familiar with exam questions beforehand, the test results lose credibility. Similarly, for AI model evaluations, it has been observed that benchmarks released before the creation date of LLM’s training data generally perform better than those released afterward [90]. Increasingly, the AI evaluation is being questioned regarding data contamination [123]. Therefore, controlling question exposure rates is a necessary measure to improve the reliability of assessments.
7.2 Fairness
@@
Fairness is a topic of profound societal significance in both education and machine learning research fields, sparking numerous discussions and leading to the development of many fairness-aware learning algorithms [35, 20, 94]. As a technology with potential applications in high-stakes testing, fairness in CAT is a paramount concern. In CAT, the bias that leads to fairness issues can be introduced through three components:
@@
  • Bias in Measurement Models. Biases in measurement models may stem from the skewed training data, which could reflect the underrepresentation of certain groups or pre-existing educational disparities [100, 152, 82]. Such biases can lead to an inaccurate and biased estimation of an examinee’s proficiency $\theta$, resulting in unfair outcomes. A practical mitigation is to evaluate calibration and model fit across subpopulations (e.g., invariance checks) and apply multi-group calibration when needed. Fairness-aware calibration objectives can also be used as a light regularizer to reduce spurious group effects.
@@
  • Bias in Question Bank. The question bank may contain biases if questions are not equally applicable or relatable to all examinees, potentially disadvantaging certain groups [19, 65]. For example, some questions in NAPLAN have been deemed unfair for rural examinees, as these questions don’t relate to their real-life experiences [131]. Various methods have been proposed to detect this type of bias [34, 19, 113]. For example, mitigation often follows an audit–repair loop: DIF analyses flag potentially biased items, which are then revised, replaced, or retired. This is usually paired with expert review to separate unintended context bias from construct-relevant differences.
@@
  • Bias in Selection Algorithms. Selection algorithm can introduce bias since every algorithm has its own “selection preferences”. For example, the Maximum Fisher Information tends to select questions with high discrimination [106]. If such questions unexpectedly correlate with specialized knowledge known only to a specific group, bias may ensue.
@@
Concerns about fairness in CAT also stem from the fact that examinees answer different questions [62]. Equating, a technique used to ensure score equivalence across different tests, is commonly employed to address such concerns [18]. Many further studies about equating scores have been conducted [159, 78, 139]. In real-world tests such as the GRE, equating has been used to standardize scores and percentiles, taking into account the difficulty of the questions answered. This process ensures that scores can be compared fairly across different examinees worldwide. In practice, equating is complemented by routine drift checks and periodic DIF re-audits, especially when new items are added or rotated. This helps preserve comparability as the bank evolves.
7.3 Robustness
Noise in CAT can impact the precision of the estimated proficiency of an examinee, leading to potential errors in score interpretation. In CAT, noise usually refers to the random variability or measurement error that can affect the accuracy of estimation. It can arise from various sources such as test administration conditions, examinee behavior, or question characteristics. For example, an examinee may be distracted by environmental noise during the test, leading to an incorrect response that does not reflect their true ability. Alternatively, a poorly worded or ambiguous question may confuse examinees, introducing unintended variability in responses.
@@ 7.3 Robustness
To mitigate the effects of noise in CAT, a robustness factor is introduced to help stabilize the estimation of proficiency by incorporating additional information, thereby counteracting the impact of noise and improving the reliability [7]. In machine learning, various robustness techniques are employed to enhance the performance of models in the presence of noise, such as regularization methods [91], data augmentation [171], adversarial methods [45], ensemble methods [80]. In the CAT testing process, significant sources of noise such as guess and slip factors made by examinees, introduce uncertainty. For example, an examinee’s proficiency level may not be uniquely determined by their responses, as they may solve a particular question correctly using different knowledge concepts or even by guessing. The presence of noise and uncertainty poses a significant challenge to the robustness of CAT systems. Veldkamp et al. [161] consider the uncertainty in question parameters during the selection process. More recently, ensemble learning has been explored to combine multiple potential estimates at each step, thereby enhancing proficiency estimation [188].
7.4 Search Efficiency
In large-scale educational testing, efficient question selection is a critical challenge. Traditional selection algorithms often evaluate all candidate questions in a brute-force manner, resulting in a linear time complexity of $O(|\mathcal{Q}|)$, where $\mathcal{Q}$ is the question bank. This becomes a computational bottleneck in intelligent testing systems. To mitigate this, some organizations like GMAT [137] rely on manual filtering rules crafted by experts, which is labor-intensive and lacks scalability. Recent research has explored two main directions to improve efficiency:
  • Heuristic Search via PSO: Particle Swarm Optimization (PSO) has been applied in IRT-based adaptive testing [76, 87]. PSO enables parallel exploration of the search space, where each particle represents a candidate question. This parallelism accelerates convergence toward optimal selections, reducing computational burden.
  • Tree-Based Indexing: Inspired by recommendation systems and information retrieval, efficient search structures such as balanced trees have been adopted [186, 8]. Hong et al. [70] propose a Search-Efficient CAT framework that employs examinee-aware space partitioning to construct a tree-based index. This method significantly narrows the search space and avoids redundant computations across testing rounds, reducing the search complexity from $O(|\mathcal{Q}|)$ to $O(\log|\mathcal{Q}|)$.
@@ 7.4 Search Efficiency
Discussion: While accuracy and efficiency stand as primary objectives, these factors hold significant importance for practical settings, especially in high-stakes testing scenarios (e.g., competitive or selective examinations). However, consideration of these factors may inevitably reduce accuracy. For example, when considering the additional fairness to ensure equity among different groups, it might be necessary to deviate from the optimal trajectory of a well-trained selection algorithm. Thus, CAT poses a multidimensional decision-making challenge, necessitating the consideration of various factors at the same time using diverse machine learning techniques. In the Appendix, we show the underlying causes and advantages of different factors in CAT test control.
8 Evaluation
@@
Various metrics have been developed to assess the performance of CAT methods, such as correlation coefficients, bias, and measurement error [37, 154]. This section introduces two of the most extensively utilized evaluation methods: simulation of proficiency estimation and examinee score prediction.
@@
Simulation of Proficiency Estimation.The simulation of ability estimation is a foundational evaluation technique in CAT [163]. Since true proficiency ($\theta_{0}$) is unobservable, we simulate it by sampling a set of values $\{\theta_{0}^{1},\theta_{0}^{2},...,\theta_{0}^{N}\}$ to represent a virtual group of examinees. This approach enables us to further emulate the interactions between examinees (with these proficiencies) and any question from the question bank, utilizing measurement models. Consequently, the estimated final proficiency values $\hat{\theta}^{T}$ can be directly compared with the true values $\theta_{0}$. For example, by computing the Mean Square Error (MSE), i.e., $\mathbb{E}\|\hat{\theta}^{T}-\theta_{0}\|$, to evaluate the accuracy of the CAT system [16, 32].
@@
Examinee Score Prediction. In machine learning–based CAT systems, proficiency estimates are often validated by predicting whether examinees will answer unseen questions correctly. Typically, examinees are split into training, validation, and test sets (e.g., 70%-20%-10%), ensuring no overlap. The training set is used to calibrate item parameters (Section 6.1) and train selection algorithms (Sections 5.3, 5.4). During validation or testing, the responses of each examinee are further divided into a candidate set $\mathcal{Q}_{i}$ for selecting questions and a held-out meta set $\mathcal{M}_{i}$ for evaluation. The candidate set $\mathcal{Q}_{i}$ (with corresponding response label $y$) is used to simulate the CAT procedure: Selecting questions from $\mathcal{Q}_{i}$, updating proficiency estimates after each step, and then accessing estimate’s precision by predicting responses on $\mathcal{M}_{i}$. The assumption is that better score predictions reflect more accurate proficiency estimates. Thus, the binary classification metrics can be used for evaluations, e.g., Prediction Accuracy (ACC) and Area Under ROC Curve (AUC) [17].
Datasets. To evaluate the effectiveness and generalizability of a CAT system, it is crucial to use diverse datasets that not only challenge the algorithm but also reflect real-world testing scenarios. Such datasets typically contain the question bank, examinee response data, and relevant contextual data. Each of these components is essential for validating the CAT system itself in realistic settings. Three types of data can be used for this evaluation:
@@
(1) Human Educational Data: This category includes data collected from educational environments in practice, such as schools, universities, and online learning platforms. It provides insights into how examinees interact with educational content and assessments in a natural setting. The data may encompass examinee information, performance responses, learning behaviors, question characteristics, etc. We have open-sourced a comprehensive education-related dataset library: https://github.com/bigdata-ustc/EduData. It includes a range of publicly available datasets along with previously private datasets, e.g., ASSISTments [51], Junyi [21], EdNet [33], and Eedi2020 [172]. Additionally, we have provided a detailed data analysis to support further CAT research and application in educational settings, which can be found at the EduData GitHub link..
(2) AI Model Response Data: The CAT paradigm is playing a crucial role in the evaluation of AI models. In particular, proficiency estimates are used to assess performance and rank models, especially for contemporary LLM evaluations. Various large-scale benchmarks and their corresponding response data can be utilized to build and test CAT systems, such as Google’s BIG-bench [145], HuggingFace’s Open LLM Leaderboard [13], HELM [99], and AlpacaEval [98]. These benchmarks encompass a wide range of tasks, with topics spanning linguistics, mathematics, medicine, common-sense reasoning, biology, physics, social bias, programming, and beyond.
@@
(3) Simulated Datasets. These are artificially created datasets that mimic the characteristics of real examinee responses as illustrated above. They can be tailored to include specific patterns, noise levels, and distributions, allowing for controlled testing of the CAT system under various scenarios. Monte Carlo simulations can also be used to generate datasets with known properties and ground truth [154]. These datasets are useful for validating the CAT system’s capability to estimate proficiencies accurately and to adapt to the simulated changes during the testing process.
9 Opportunities for Future Research